# Constraining the redshift of PG 1553+113 using the photohadronic model

### v1

@@ Abstract
Between 2005 and 2015, the BL Lacertae object PG 1553+113 exhibited multiple very high-energy (VHE; $>100$ GeV) gamma-ray flares, which were detected by the Cherenkov telescopes, High Energy Stereoscopic System (HESS), Major Atmospheric Gamma Imaging Cherenkov (MAGIC), and the Very Energetic Radiation Imaging Telescope Array System (VERITAS). Despite the uncertainty surrounding its redshift ($z$), various studies have sought to estimate this value. In this study, seventeen independently observed VHE gamma-ray spectra of PG 1553+113 are analyzed using four distinct extragalactic background light (EBL) models alongside the photohadronic framework. A global $\chi^{2}$ fit is applied to all observational data to determine the best-fitting redshift for each EBL model. Additionally, confidence level (CL) intervals for the redshift are calculated across all EBL models. The findings demonstrate that the photohadronic framework effectively describes all observed spectra. Among the EBL models, the lowest total $\chi^{2}$ value of 58.82 (with 58 degrees of freedom) was obtained using the model from Saldana-Lopez et al. (2021), while the highest value of 75.67 was associated with the model from Dominguez et al. (2011). The 95 per cent CL intervals for the statistical error of the redshift of PG 1553+113 are $0.5370.500<z<$ 0.537 for the Saldana-Lopez et al. (2021) model and $0.5270.491<z<$ 0.527 for the Dominguez et al. (2011) model. These results highlight the consistency of the photohadronic approach in interpreting the observed VHE gamma-ray spectra.
@@ 1 INTRODUCTION
In BL Lac objects, the non-thermal radiation significantly overshadows the stellar light emitted by their host galaxies, making it challenging to determine accurately the redshifts. This limitation creates uncertainties in studying the cosmic evolution, deciphering the source’s characteristics, and analyzing its intrinsic VHE spectrum. Given that VHE photons are absorbed by the EBL, obtaining precise redshift measurements is essential for understanding the EBL’s influence on VHE photons.
PG 1553+113, a well-known blazar, was first identified in a photographic survey (Green et al., 1986). Its spectral properties have been thoroughly examined, leading to its classification as a high-frequency synchrotron peaked BL Lac (HBL) (Chang et al., 2019). Over the years, multiple VHE observations have been carried out using different instruments. The HESS telescope array first detected this blazar in gamma-rays in 2005 (Aharonian et al., 2006), and it was later monitored by MAGIC and VERITAS in various campaigns between 2005 and 2015, especially during periods of heightened activity.
@@
where $E_{\gamma}$ denotes the energy of the observed VHE photon, $F_{\gamma}$ is the observed flux, and $F_{\rm int}$ represents the intrinsic flux. The optical depth, $\tau_{\gamma\gamma}$, which arises from the pair production process $\gamma\gamma\rightarrow e^{+}e^{-}$, is a function of both $E_{\gamma}$ and the redshift $z$ of the source. The exponential term in Eq. (1) signifies the reduction in VHE flux caused by $e^{+}e^{-}$ pair production (Stecker et al., 1992; Ackermann et al., 2012; Padovani et al., 2017). As a result, determining the redshift with precision is critical for deriving the intrinsic flux from the observed measurements. To analyze VHE gamma-ray spectra from sources at different redshifts, the Imaging Atmospheric Cherenkov Telescopes (IACTs) collaborations rely on well-established EBL models developed by Franceschini et al. (2008); Finke et al. (2010); Dominguez et al. (2011); Gilmore et al. (2012); Saldana-Lopez et al. (2021).
The redshift of PG 1553+113 has not been directly measured but has been inferred through various methods. Using data from HESS observations in 2005 and considering the absorption of VHE gamma-rays by the EBL, Aharonian et al. (2006) derived an upper limit of $z<0.74$. Later, Danforth et al. (2010) established a lower limit of $0.4<z$ by detecting intervening Ly $\alpha$ absorbers, while also setting an upper limit of $z<0.58$ due to the absence of Ly $\beta$ absorbers at $z>0.4$. Subsequent studies have supported these constraints. For instance, Aliu et al. (2015) analyzed the time-averaged VHE spectrum from VERITAS observations between 2010 and 2012, incorporating EBL evolution, and proposed a redshift upper limit of $z\leq 0.62$. More recently, Dorigo Jones et al. (2022) used a Ly $\alpha$-forest-based approach to suggest a narrower redshift range of $0.4360.408<z<$ 0.436. Our study aims to estimate the redshift of PG 1553+113 using the photohadronic model, employing a statistical approach and a comprehensive dataset.
The photohadronic model (Sahu, 2019) has previously been applied successfully to interpret the VHE gamma-ray spectra of HBLs and extreme HBLs (EHBLs) (Sahu et al., 2020, 2021, 2022). In this work, the photohadronic model (Sahu, 2019) is combined with four established EBL models (Franceschini et al. (2008); Dominguez et al. (2011); Gilmore et al. (2012); Saldana-Lopez et al. (2021)) to analyze seventeen VHE flaring events of PG 1553+113 observed between 2005 and 2015 by the HESS, VERITAS, and MAGIC collaborations. By performing a global $\chi^{2}$ fit to the experimental data from these VHE spectra, central values and CL intervals for the redshift of PG 1553+113 are determined for each EBL model. This method has been validated in prior analyses, such as the redshift constraint determination for the BL Lac object PKS 1424+240, which aligned with spectroscopic measurements (Sahu et al., 2024).
@@ 2 VHE Observations of PG 1553+113
This section provides a concise overview of the seventeen observations of PG 1553+113 by different Cherenkov telescopes which we analyze here to constrain the redshift. Key details for each observation are summarized in Table 1. To simplify referencing, each observation is labeled according to the telescope and the year it was conducted. For instance, the HESS observation of 2005 is labeled as H05.
@@ 2.1 H05, H06
HESS conducted VHE observations of PG 1553+113 in 2005 and 2006. Near-infrared spectroscopy in the $H+K$ band (1.45-2.45 $\mu$m) was performed in March 2006 using the Spectrograph for INtegral Field Observations in the Near Infrared (SINFONI), an integral field spectrometer on the European Southern Observatory (ESO) Very Large Telescope (VLT) in Chile.
Over the two years of observations (totaling 24.8 hours of live time), HESS detected a VHE signal with a significance of approximately 10 standard deviations. No significant flux or spectral variations were observed on any time scale within the VHE data. Despite the high sensitivity of the SINFONI spectrum, which is among the most detailed near-infrared spectra ever obtained for this object, no absorption or emission lines were identified, preventing a determination of the redshift.
A total of 30.3 hours of HESS observations were conducted, typically in segments of about 28 minutes (runs), during 2005 and 2006. The telescopes were offset by ±$\pm 0.5^{\circ}$ from the position of PG 1553+113 during these observations. After applying data-quality selection criteria, the useful exposure time amounted to 24.8 hours of live time, with 9 of the 66 runs excluded. The 2005 observation, referred to as H05, had a live time of 7.6 hours, while the 2006 observation, labeled H06, had a live time of 17.2 hours.
@@ 2.2 H12
As discussed previously, the observations H05 and H06 showed no signs of variability and are collectively termed as the "pre-flare" data set. This pre-flare data set comprises 26.4 hours of high-quality live time observations. During the 2012 flaring period (referred here as H12), eight observational runs, each lasting approximately 28 minutes were conducted on April 26 and 27, totaling 3.5 hours of live time. All observations were performed in wobble mode, where the source is positioned at an offset of 0.5 from the centre of the instrument’s field of view. This method resulted in acceptance-corrected live times of 24.8 hours for the pre-flare data and 3.2 hours for the flare data.
@@ 2.3 H13
The HESS phase II, with a fifth telescope CT5, observed the HBL PG 1553+113 in 2013 between May 29 and August 9, 2013 (MJD 56441–56513) (H.E.S.S. Collab. et al., 2017), referred to as H13 in this work. These observations are divided into 28-minute segments called runs and conducted in wobble mode, where the camera’s field of view is offset by $0.5^{\circ}$ or $0.7^{\circ}$ from the source position, either along the right-ascension or declination axis. For this analysis, only runs with the source positioned between $0.35^{\circ}$ and $1.2^{\circ}$ from the camera centre were included. During the commissioning phase, some runs employed non-standard wobble offsets to evaluate the instrument’s performance. This approach ensures the source remains well within the field of view and facilitates accurate background subtraction.
@@ 2.5 M06
The MAGIC telescope observed PG 1553+113 during a multiwavelength campaign from July 14 to July 27, 2006, accumulating 9.5 hours of data at zenith angles ranging between $18^{\circ}$ and $35^{\circ}$ (Albert et al., 2009). These observations are labeled as M06 in this study. The data were collected in wobble mode, with the source positioned at an offset of ±$\pm 0.4^{\circ}$ from the camera’s centre, allowing simultaneous measurement of both the source and the background. Approximately one hour of data was discarded due to technical issues.
@@ 2.6 M07, M08, M09
Following its initial detection, PG 1553+113 became a regular target for observations with the MAGIC telescope. As reported in reference Aleksić et al. (2012), new data collected in 2007, 2008, and 2009 were analyzed, and here we refer to these observations as M07, M08, and M09, respectively. Since the autumn of 2009, MAGIC has operated as a stereo system consisting of two IACTs located on La Palma in the Canary Islands, Spain. However, the data discussed here were acquired before this upgrade, using only the single MAGIC I telescope.
As mentiones previously, PG 1553+113 was observed by MAGIC for approximately 19 hours in 2005 and 2006 (Albert et al., 2007) and was also the focus of a multiwavelength campaign in July 2006 involving optical, X-ray, and TeV gamma-ray  (Albert et al., 2009). Reference Aleksić et al. (2012) presents the results of follow-up observations conducted for 14 hours in March–April 2007, nearly 26 hours in March–May 2008 (some of which were simultaneous with observations by other instruments (Aleksić et al., 2010)), and about 24 hours in March–July 2009, with some of the latter taken under moderate moonlight conditions. Unfortunately, the 2008 and 2009 observations were significantly impacted by adverse weather, including calima (Saharan sand dust in the atmosphere), which reduced the usable data and increased the energy threshold.
@@ 2.7 M08II
A comprehensive simultaneous multi-frequency observational campaign was conducted between March and April 2008, involving data collection across optical, X-ray, high-energy (HE) gamma-ray, and VHE gamma-ray bands (Aleksić et al., 2010). The average VHE differential spectrum of PG 1553+113 derived from these observations is reported and labeled as M08II in this work.
@@ 2.9 M12II, M12III, M13, M14, M15
Among these, five data sets correspond to the BL Lac object PG 1553+113, covering observations from 2012 to 2015 with a combined observation time of 66.36 hours. These data sets, as reported in reference Acciari et al. (2019), are labeled as PG1553+113_0202 (referred to here as M12II), PG1553+113_0203 (M12III), PG1553+113_0302 (M13), PG1553+113_0303 (M14), and PG1553+113_0306 (M15).
@@ 2.10 V1012
Observations of PG 1553+113 were conducted by VERITAS from May 2010 to June 2012, totaling 95 hours of observation time. Data were collected in wobble mode, with the source positioned at a 0.5-degree offset from the telescope’s pointing direction to enable simultaneous background estimation. The zenith angles for these observations ranged from 20 to 30 degrees, with an average of 23 degrees. The combination of low zenith angles and event selection cuts optimized for soft-spectrum sources resulted in an analysis energy threshold (the energy at which the photon rate peaks after cuts) of 180 GeV.
The analysis used a circular signal region centered on the source’s nominal position, extending radially outward by 0.14 degrees. After applying quality selection criteria based on weather conditions and instrument stability, and accounting for instrument read-out dead time, a total of 80 hours of live time were retained. These observations yielded an overall detection significance of 53 standard deviations, with the excess consistent with a point-like source at the location of PG 1553+113.
@@ 3 THE PHOTOHADRONIC MODEL
where $\tau_{\gamma\gamma}(E_{\gamma},z)$ accounts for the absorption of gamma-rays due to interactions with the EBL. The spectral index $\delta$ is defined as $\delta=\alpha+\beta$, where $\alpha$ is the proton spectral index and $\beta$ represents the power-law behavior of the seed photon spectrum in the low-energy tail of the SSC spectrum. The normalization factor $F_{0}$ is determined from observational data, while $\delta$ serves as the sole free parameter in the photohadronic model (Sahu et al., 2019). As evident from Eq. (6), any curvature in the spectrum can be attributed to the exponential absorption term (Acciari et al., 2019).
@@ 4 ANALYSIS AND RESULTS
We analyze all the seventeen observations discussed in Sec. 2 in the context of the photohadronic model to constrain the redshift of the HBL PG 1553+113. To account for the EBL correction to the VHE gamma-ray spectra, we use four well known and widely used EBL models. Here, these models are referred to as Saldana (Saldana-Lopez et al., 2021), Franceschini (Franceschini et al., 2008), Domínguez (Dominguez et al., 2011), and Gilmore (Gilmore et al., 2012).
@@
To analyze the VHE spectra and determine the best fits, each EBL model is integrated into the photohadronic framework. A global $\chi^{2}$ fit is performed on all data points by simultaneously varying the redshift $z$ (common to all observations), the normalization constants $F_{0}$, and the spectral parameter $\delta$ for each of the seventeen independent observations. This global fitting procedure is repeated for all four EBL models to identify the best-fitting values of $z$, $F_{0}$, and $\delta$. Confidence level intervals for the statistical errors of the redshift at 68, 90, and 95 per cent of coverage probability are then calculated for each EBL model.

### v2

@@ Abstract
Between 2005 and 2015, the BL Lacertae object PG 1553+113 exhibited multiple very high-energy (VHE; $>100$ GeV) gamma-ray flares, which were detected by the Cherenkov telescopes, High Energy Stereoscopic System (HESS), Major Atmospheric Gamma Imaging Cherenkov (MAGIC), and the Very Energetic Radiation Imaging Telescope Array System (VERITAS). Despite the uncertainty surrounding its redshift ($z$), various studies have sought to estimate this value. In this study, seventeen independently observed VHE gamma-ray spectra of PG 1553+113 are analysed using four distinct extragalactic background light (EBL) models alongside the photohadronic framework. A global $\chi^{2}$ fit is applied to all observational data to determine the best-fitting redshift for each EBL model. Additionally, confidence level (CL) intervals for the redshift are calculated across all EBL models. The findings demonstrate that the photohadronic framework effectively describes all observed spectra. Among the EBL models, the lowest total $\chi^{2}$ value of 58.82 (with 58 degrees of freedom) was obtained using the model from Saldana-Lopez et al. (2021), while the highest value of 75.67 was associated with the model from Domínguez et al. (2011). The 95 per cent CL intervals for the statistical error of the redshift of PG 1553+113 are $0.5370.500<z<$ 0.537 for the Saldana-Lopez model and $0.5270.491<z<$ 0.527 for the Domínguez model. These results highlight the consistency of the photohadronic approach in interpreting the observed VHE gamma-ray spectra.
@@ 1 INTRODUCTION
In BL Lac objects, the non-thermal radiation significantly overshadows the stellar light emitted by their host galaxies, making it challenging to determine accurately the redshifts. This limitation creates uncertainties in studying the cosmic evolution, deciphering the source’s characteristics, and analysing its intrinsic VHE spectrum. Given that VHE photons are absorbed by the EBL, obtaining precise redshift measurements is essential for understanding the EBL’s influence on VHE photons.
PG 1553+113, a well-known blazar, was first identified in a photographic survey (Green et al., 1986). Its spectral properties have been thoroughly examined, leading to its classification as a high-frequency synchrotron peaked BL Lac (HBL) (Chang et al., 2019). Over the years, multiple VHE observations have been carried out using different instruments. The HESS telescope array first detected this blazar in gamma-rays in 2005 (Aharonian et al., 2006), and it was later monitored by MAGIC and VERITAS in various campaigns between 2005 and 2015, especially during periods of heightened activity. PG 1553+113 is a prime target for high-energy astrophysics due to its extreme variability and emission mechanisms. However, its redshift remains unconfirmed via spectroscopy, making indirect estimation methods critical for interpreting its emission properties and cosmological impact.
@@
where $E_{\gamma}$ denotes the energy of the observed VHE photon, $F_{\gamma}$ is the observed flux, and $F_{\rm int}$ represents the intrinsic flux. The optical depth, $\tau_{\gamma\gamma}$, which arises from the pair production process $\gamma\gamma\rightarrow e^{+}e^{-}$, is a function of both $E_{\gamma}$ and the redshift $z$ of the source. The exponential term in Eq. (1) signifies the reduction in VHE flux caused by $e^{+}e^{-}$ pair production (Stecker et al., 1992; Ackermann et al., 2012; Padovani et al., 2017). As a result, determining the redshift with precision is critical for deriving the intrinsic flux from the observed measurements. To analyse VHE gamma-ray spectra from sources at different redshifts, the Imaging Atmospheric Cherenkov Telescopes (IACTs) collaborations rely on well-established EBL models developed by Franceschini et al. (2008); Finke et al. (2010); Domínguez et al. (2011); Gilmore et al. (2012); Saldana-Lopez et al. (2021).
PG 1553+113 lacks a direct spectroscopic redshift due to its featureless optical spectrum (a hallmark of BL Lacs) but has been inferred through various methods. Using data from HESS observations in 2005 and considering the absorption of VHE gamma-rays by the EBL, Aharonian et al. (2006) derived an upper limit of $z<0.74$. Later, Danforth et al. (2010) established a lower limit of $0.4<z$ by detecting intervening Ly $\alpha$ absorbers, while also setting an upper limit of $z<0.58$ due to the absence of Ly $\beta$ absorbers at $z>0.4$. Subsequent studies have supported these constraints. For instance, Aliu et al. (2015) analysed the time-averaged VHE spectrum from VERITAS observations between 2010 and 2012, incorporating EBL evolution, and proposed a redshift upper limit of $z\leq 0.62$. More recently, Dorigo Jones et al. (2022) used a Ly $\alpha$-forest-based approach to suggest a narrower redshift range of $0.4360.408<z<$ 0.436. Our study aims to estimate the redshift of PG 1553+113 using the photohadronic model, employing a statistical approach and a comprehensive dataset.
The photohadronic model (Sahu, 2019) has previously been applied successfully to interpret the VHE gamma-ray spectra of HBLs and extreme HBLs (EHBLs) (Sahu et al., 2020, 2021, 2022). In this work, the photohadronic model (Sahu, 2019) is combined with four established EBL models (Franceschini et al. (2008); Domínguez et al. (2011); Gilmore et al. (2012); Saldana-Lopez et al. (2021)) to analyse seventeen VHE flaring events of PG 1553+113 observed between 2005 and 2015 by the HESS, VERITAS, and MAGIC collaborations. By performing a global $\chi^{2}$ fit to the experimental data from these VHE spectra, central values and CL intervals for the redshift of PG 1553+113 are determined for each EBL model. This method has been validated in prior analyses, such as the redshift constraint determination for the BL Lac object PKS 1424+240, which aligned with spectroscopic measurements (Sahu et al., 2024).
@@ 2 VHE Observations of PG 1553+113
This section provides a concise overview of the seventeen observations of blazar PG 1553+113 analysed in this study to constrain its redshift. The data were collected by three major Cherenkov telescopes: MAGIC (Moralejo & MAGIC Collaboration, 2004), HESS (Hinton & HESS Collaboration, 2004), and VERITAS (Holder et al., 2006). Key details for each observation are summarized in Table 1. To simplify referencing, each observation is labelled according to the telescope and the year it was conducted. For instance, the HESS observation of 2005 is labeled as H05.
@@ 2.1 H05, H06
The HESS system observed PG 1553+113 in the VHE gamma-ray regime during observational campaigns in 2005 and 2006. Over the two-year period, HESS accumulated 24.8 hours of live time, resulting in the detection of a VHE gamma-ray signal with a statistical significance of approximately 10 standard deviations. The emission displayed remarkable stability, with no significant flux variability or spectral changes observed across any timescale.
Complementary near-infrared spectroscopic observations targeting the $H+K$ band (1.45–2.45 $\mu$m) were conducted in March 2006 using the Spectrograph for INtegral Field Observations in the Near Infrared (SINFONI) on the European Southern Observatory’s Very Large Telescope (VLT) in Chile. Despite the exceptional sensitivity of the SINFONI integral field spectrometer, producing one of the highest-resolution near-infrared spectra of PG 1553+113 to date, the data revealed no detectable absorption or emission features. This absence of spectral lines precluded any direct constraints on the source’s redshift, leaving its cosmological distance uncertain.
A total of 30.3 hours of HESS observations were conducted, typically in segments of about 28 minutes (runs), during 2005 and 2006. The telescopes were offset by ±$\pm 0.5^{\circ}$ from the position of PG 1553+113 during these observations. After applying data-quality selection criteria, the useful exposure time amounted to 24.8 hours of live time, with 9 of the 66 runs excluded. The 2005 observation, referred to as H05, had a live time of 7.6 hours, while the 2006 observation, labelled H06, had a live time of 17.2 hours.
@@ 2.2 H12
As discussed previously, the observations H05 and H06 showed no signs of variability and are collectively termed as the "pre-flare" data set. This pre-flare data set comprises 26.4 hours of high-quality live time observations. During the 2012 flaring period (referred here as H12), eight observational runs, each lasting approximately 28 minutes were conducted on April 26 and 27, totalling 3.5 hours of live time. All observations were performed in wobble mode, where the source is positioned at an offset of 0.5 from the centre of the instrument’s field of view. This method resulted in acceptance-corrected live times of 24.8 hours for the pre-flare data and 3.2 hours for the flare data.
@@ 2.3 H13
The HESS phase II, with a fifth telescope CT5, observed the HBL PG 1553+113 in 2013 between May 29 and August 9, 2013 (MJD 56441–56513) (H.E.S.S. Collab. et al., 2017), referred to as H13 in this work. These observations are divided into 28-minute segments called runs and conducted in wobble mode, where the camera’s field of view is offset by $0.5^{\circ}$ or $0.7^{\circ}$ from the source position, either along the right-ascension or declination axis. For this analysis, only runs with the source positioned between $0.35^{\circ}$ and $1.2^{\circ}$ from the camera centre were included. During the commissioning phase, some runs employed non-standard wobble offsets to evaluate the instrument’s performance. This approach ensures the source remains well within the field of view and facilitates accurate background subtraction. The analysis comprises 39 runs with 16.8 hours of live time.
@@ 2.5 M06
The MAGIC telescope observed PG 1553+113 during a multiwavelength campaign from July 14 to July 27, 2006, accumulating 9.5 hours of data at zenith angles ranging between $18^{\circ}$ and $35^{\circ}$ (Albert et al., 2009). These observations are labelled as M06 in this study. The data were collected in wobble mode, with the source positioned at an offset of ±$\pm 0.4^{\circ}$ from the camera’s centre, allowing simultaneous measurement of both the source and the background. Approximately one hour of data was discarded due to technical issues.
@@ 2.6 M07, M08, M09
Following its initial detection, PG 1553+113 became a regular target for observations with the MAGIC telescope. As reported in reference Aleksić et al. (2012), new data collected in 2007, 2008, and 2009 were analysed, and here we refer to these observations as M07, M08, and M09, respectively. Since the autumn of 2009, MAGIC has operated as a stereo system consisting of two IACTs located on La Palma in the Canary Islands, Spain. However, the data discussed here were acquired before this upgrade, using only the single MAGIC I telescope.
As mentioned previously, PG 1553+113 was observed by MAGIC for approximately 19 hours in 2005 and 2006 (Albert et al., 2007) and was also the focus of a multiwavelength campaign in July 2006 involving optical, X-ray, and TeV gamma-ray  (Albert et al., 2009). Reference Aleksić et al. (2012) presents the results of follow-up observations conducted for 14 hours in March–April 2007, nearly 26 hours in March–May 2008 (some of which were simultaneous with observations by other instruments (Aleksić et al., 2010)), and about 24 hours in March–July 2009, with some of the latter taken under moderate moonlight conditions. Unfortunately, the 2008 and 2009 observations were significantly impacted by adverse weather, including calima (Saharan sand dust in the atmosphere), which reduced the usable data and increased the energy threshold.
@@ 2.7 M08II
A comprehensive simultaneous multi-frequency observational campaign was conducted between March and April 2008, involving data collection across optical, X-ray, high-energy (HE) gamma-ray, and VHE gamma-ray bands (Aleksić et al., 2010). The average VHE differential spectrum of PG 1553+113 derived from these observations is reported and labelled as M08II in this work.
@@ 2.9 M12II, M12III, M13, M14, M15
Among these, five data sets correspond to the BL Lac object PG 1553+113, covering observations from 2012 to 2015 with a combined observation time of 66.36 hours. These data sets, as reported in reference Acciari et al. (2019), are labelled as PG1553+113_0202 (referred to here as M12II), PG1553+113_0203 (M12III), PG1553+113_0302 (M13), PG1553+113_0303 (M14), and PG1553+113_0306 (M15).
@@ 2.10 V1012
Observations of PG 1553+113 were conducted by VERITAS from May 2010 to June 2012, totalling 95 hours of observation time. Data were collected in wobble mode, with the source positioned at a 0.5-degree offset from the telescope’s pointing direction to enable simultaneous background estimation. The zenith angles for these observations ranged from 20 to 30 degrees, with an average of 23 degrees. The combination of low zenith angles and event selection cuts optimized for soft-spectrum sources resulted in an analysis energy threshold (the energy at which the photon rate peaks after cuts) of 180 GeV.
The analysis used a circular signal region centred on the source’s nominal position, extending radially outward by 0.14 degrees. After applying quality selection criteria based on weather conditions and instrument stability, and accounting for instrument read-out dead time, a total of 80 hours of live time were retained. These observations yielded an overall detection significance of 53 standard deviations, with the excess consistent with a point-like source at the location of PG 1553+113.
@@ 3 THE PHOTOHADRONIC MODEL
where $\tau_{\gamma\gamma}(E_{\gamma},z)$ accounts for the absorption of gamma-rays due to interactions with the EBL. The spectral index $\delta$ is defined as $\delta=\alpha+\beta$, where $\alpha$ is the proton spectral index and $\beta$ represents the power-law behaviour of the seed photon spectrum in the low-energy tail of the SSC spectrum. The normalization factor $F_{0}$ is determined from observational data, while $\delta$ serves as the sole free parameter in the photohadronic model (Sahu et al., 2019). As evident from Eq. (6), any curvature in the spectrum can be attributed to the exponential absorption term (Acciari et al., 2019).
@@ 4 ANALYSIS AND RESULTS
We analyse all the seventeen observations discussed in Sec. 2 in the context of the photohadronic model to constrain the redshift of the HBL PG 1553+113. To account for the EBL correction to the VHE gamma-ray spectra, we use four well known and widely used EBL models. Here, these models are referred to as Saldana (Saldana-Lopez et al., 2021), Franceschini (Franceschini et al., 2008), Domínguez (Domínguez et al., 2011), and Gilmore (Gilmore et al., 2012).
@@
To analyse the VHE spectra and determine the best fits, each EBL model is integrated into the photohadronic framework. A global $\chi^{2}$ fit is performed on all data points by simultaneously varying the redshift $z$ (common to all observations), the normalization constants $F_{0}$, and the spectral parameter $\delta$ for each of the seventeen independent observations. This global fitting procedure is repeated for all four EBL models to identify the best-fitting values of $z$, $F_{0}$, and $\delta$. Confidence level intervals for the statistical errors of the redshift at 68, 90, and 95 per cent of coverage probability are then calculated for each EBL model.